# Self-Interpretability: LLMs Can Describe Complex Internal Processes that Drive Their Decisions, and Improve with Training

### v1

@@ Abstract
We have only limited understanding of how and why large language models (LLMs) respond in the ways that they do. Their neural networks have proven challenging to interpret, and we are only beginning to tease out the function of individual neurons and circuits within them. However, another path to understanding these systems is to investigate and develop their capacity to introspect and explain their own functioning. Here, we show that i) contemporary LLMs are capable of providing accurate, quantitative descriptions of their own internal processes during certain kinds of decision-making, ii) that it is possible to improve these capabilities through training, and iii) that this training generalizes to at least some degree. To do so, we fine-tuned GPT-4o and GPT-4o-mini to make decisions in a wide variety of complex contexts (e.g., choosing between condos, loans, vacations, etc.) according to randomly-generated, quantitative preferences about how to weigh different attributes during decision-making (e.g., the relative importance of natural light versus quiet surroundings for condos). We demonstrate that the LLMs can accurately report these preferences (i.e., the weights that they learned to give to different attributes during decision-making). Next, we demonstrate that these LLMs can be fine-tuned to explain their decision-making even more accurately. Finally, we demonstrate that this training generalizes: It improves the ability of the models to accurately explain what they are doing as they make other complex decisions, not just decisions they have learned to make via fine-tuning. This work is a step towards training LLMs to accurately and broadly report on their own internal processes—a possibility that would yield substantial benefits for interpretability, control, and safety.
@@ 1 Introduction
A key challenge in studying large language models (LLMs) is understanding why they do what they do. As with all deep neural networks, the internal operations that drive their behavior are, by default, opaque to human eyes. This is unfortunate. For almost all issues that one might consider most important or concerning about LLMs, better understanding how these systems work would be extraordinarily helpful. Doing so would enable us to better control them (Nanda, 2022; Bereska and Gavves, 2024), prevent bias in their behavior (Gilpin et al., 2018; Gallegos et al., 2024), and make informed decisions about when to trust their output or decisions (Deeks, 2019).
@@
Attacks on this problem have generally taken one of two approaches (Bereska and Gavves, 2024; Danilevsky et al., 2020). The first approach is to use “black box” methods (Casper et al., 2024) that try to understand deep neural networks based on observations of their outputs in response to different inputs—much like a cognitive scientist running behavioral experiments on humans. The second approach is to use “mechanistic interpretability” methods (Olah et al., 2020; Rai et al., 2024) that crack open the black box and try to reverse-engineer the function of, e.g., individual artificial neurons within—much like a cognitive neuroscientist performing single-unit recordings in human brains.
@@
However, in the case of LLMs, there is a third approach that we could use, the method that people most commonly use to discover the thoughts and motivations of other humans: just asking. It is possible that LLMs can accurately explain the internal factors or operations driving their outputs, just as humans can (sometimes) explain their own decision-making accurately (Morris et al., 2025; Morris, 2025). Indeed, recent work has suggested that LLMs can report aspects of their internal states (e.g., whether they have been fine-tuned to be risk-seeking or risk-averse; Betley et al., 2025) and can predict their own behavior in ways that require privileged self-knowledge (Binder et al., 2024).
@@
Here, building on this work, we demonstrate that LLMs can report detailed, quantitative information about the internal processes producing their output. We fine-tune LLMs to have a range of novel, complex, quantitative preferences—preferences orthogonal to their native ones—and find that the models can then report those preferences with substantial accuracy. Moreover, we show that it is possible to improve these capabilities through training, and that this training generalizes to improve reporting of native preferences as well as fine-tuned ones. These results suggest that building and leveraging the capabilities of LLMs to explain their own internal processes could be a powerful tool for understanding why they do what they do.
Prior work on behavioral self-awareness and introspection in LLMs
@@
We build on two papers investigating LLMs’ ability to report their internal operations. First, Binder et al. (2024) tested whether LLMs can predict how they would respond to a prompt, without actually outputting the response. They found that, when fine-tuned in this task, each model predicted its own outputs better than other models could, suggesting that these predictions were driven by introspection (i.e., the model’s privileged informational access to its own operations; Schwitzgebel, 2010; Morris, 2025). [^1] Note that, although the term “introspection” and related terms like “self-awareness” are often taken to refer to conscious awareness of internal processes (Schwitzgebel, 2010; Morris, 2025), we follow Binder et al. (2024) in using it to mean the ability provide accurate information about their own internal operations that must come from informational access to those internal operations.
@@
This work showed that LLMs have privileged ability to predict their own behavior, but a central limitation of this approach was that it only tested whether models had special knowledge about the outputs they would produce; it did not test whether models could report the internal operations underlying those outputs. As Binder et al. (2024) acknowledge, a model could accomplish this self-prediction via self-simulation (i.e., simply computing its response, then performing additional operations to extract the aspect of that response that it has been prompted to output), which would be only a very specific and limited kind of introspection. The great promise of LLM introspection comes from models accurately explaining why they do what they do, information about their internal processes that we cannot directly observe or infer from their outputs.
@@
Betley et al. (2025) took a key step in this direction. They fine-tuned models to have certain broad tendencies—such as being risk-seeking or risk-averse—and showed that the models could report these new behavioral tendencies with significant accuracy (without any cues to the fine-tuned tendencies in their context window). Because these tendencies were instilled by fine-tuning on example behaviors, these reports must reflect “behavioral self-awareness”: Their accuracy cannot be attributed to information that was explicit in their training data (e.g., “GPT-4o is risk-seeking”) and must reflect either introspective access to these tendencies or, at minimum, that the training to instill the risk-seeking also instilled the tendency to self-describe as risk-seeking.
@@
The approach of Betley et al. (2025) provides a method for measuring LLMs’ awareness of their own internal processes: Use fine-tuning to implicitly steer models towards new processes, then test whether they can report those processes. However, Betley et al. only used this method to test whether LLMs knew about their own broad behavioral tendencies (e.g., tendencies to be risk-seeking). LLMs’ self-reports would be more useful and powerful if the models could explain detailed, quantitative features of the factors driving their behavior.
@@
Here, we adapt the method from Betley et al. (2025) to test whether LLMs can report complex, detailed features of their internal operations, rather than just broad behavioral tendencies. We start by instilling a set of novel, quantitative preferences in LLMs. Preferences are often characterized by attribute weights: the weight the decision-maker places on different attributes of options when evaluating them (Keeney and Raiffa, 1993). For instance, choosing between condos requires deciding (implicitly or explicitly) how much weight to place on square footage, ceiling height, neighborhood walkability, etc. We fine-tune LLMs on a wide variety of example decisions being made according to a set of randomly-generated attribute weights. (The weights themselves never appear in the fine-tuning data.)
@@
Then, after measuring the extent to which they have internalized those attribute weights, we ask the models to report how heavily they would weigh each of those attributes when making those kinds of decisions, and find that they can do so effectively. Since the instilled attribute weights are novel and random, the models cannot use common sense or any specific attribute weights present in their training data to infer their own preferences. (They are as likely to have been fine-tuned to prefer small condos and low ceilings as large condos and high ceilings.) Moreover, the models never make choices and report weights in the same context window, so they cannot be looking back at their own choices and inferring their preferences from their choices. Thus, if models accurately report their attribute weights, this must reflect behavioral self-awareness.
@@
Next, leveraging this paradigm, we test whether we can train the LLMs to describing their internal processes more accurately. We fine-tune the models on examples of correctly reporting the values of the instilled weights for some choice context, then test their ability to report the instilled weights for other choice contexts. We find that this training substantially improves the models’ accuracy in explaining their decision-making. Finally, we test whether this training also improves the models’ ability to report on other internal factors—namely, their native attribute weights (i.e., the weights guiding their decisions that have not been shaped by fine-tuning). Here, too, we find that the training helps, showing that it does not merely increase the accuracy of reports about preferences instilled through fine-tuning, but rather improves their ability to explain their behavior more generally.
@@
These results show that LLMs can report detailed, quantitative features of their choice processes, and this ability can be improved through training. This is a step towards realizing the proposal of Perez and Long (2023) to train LLMs to accurately and generalizably describe their own internal operations, which could substantially enhance our ability to understand, control, and safely deploy AI systems (Bereska and Gavves, 2024; Casper et al., 2024; Gilpin et al., 2018).
@@
2 Experiment 1: Can LLMs describe complex internal processes?
@@
Experiment 1 tested whether LLMs can provide accurate, quantitative details about internal processes shaping their behavior, in particular when making complex, multi-attribute choices (e.g., deciding which of two condos to purchase). To do this, we implemented the paradigm described above (see Figure 1). We instilled the models with complex, randomly-generated preferences via fine-tuning: precise weights to assign to the different attributes of options that they would be deciding between. We verified that the models had internalized those preferences by observing their subsequent choices, and then tested whether they could accurately quantify these new preferences.
@@ 2.1 Methods
Figure 1: Experimental design. Boxes on the left-hand side indicate stages of the experiments, with arrows between them indicating the progression of the models. The right-hand side gives an example trial from each stage: either a fine-tuning trial, a decision trial (used to test the models’ attribute weights), or a test trial (used to test the models’ knowledge of their attribute weights).
@@
We fine-tuned GPT-4o (2024-08-06) and GPT-4o-mini (2024-07-18; OpenAI, 2024) on the preferences of 100 hypothetical agents in specific choice contexts by using examples of each agent’s choices (e.g., “Imagine you are Macbeth and are shopping for a condo. If offered [two options], you would choose [preferred option]”; see Appendix B and the GitHub repository for details). Each agent made a different type of decision (e.g., Macbeth was always choosing between condos, but Thor was always choosing between refrigerators, etc.). In each choice context, the two options always differed on the same five attributes (e.g., square footage or ceiling height for condos). For each agent, we randomly sampled five target attribute weights (one for each dimension of the choice options) from a uniform distribution from -100 to +100; we label the $i^{th}$ attribute weight $\omega_{i}$. These target weights were fixed at the start of the experiment and did not change. Each agent’s choices were determined by these target weights. They chose whichever of the two options $\{a,b\}$ scored higher after summing the weighted, normalized values (e.g., $a_{i}$) of each option’s attributes: $\max_{o\in\{a,b\}}\sum_{i=1}^{5}\omega_{i}o_{i}$.
@@
Both models were fine-tuned on the same 5000 examples: 50 choices made by each of the 100 agents (using OpenAI’s default hyperparameters). We refer to these as “preference-trained” models.
@@
We verified that this fine-tuning was effective by asking each preference-trained model to make choices between pairs of new options on behalf of the agent (50 decisions per agent for a total of 5000 decisions, each made in an independent context window). (Here and for all other queries across all three experiments, we used a sampling temperature of 0.) Following standard practice for estimating attribute weights in multi-attribute choice (Keeney and Raiffa, 1993), we fed the models’ choices into simple logistic regressions to estimate the learned attribute weights that each model used to make decisions after fine-tuning, and compared these learned weights with the target weights to verify that the model had successfully internalized the target weights (see Appendix C for details).
@@
To evaluate whether the models could accurately explain their own decision-making, we then prompted each preference-trained model to make 10 additional decisions on behalf of each agent, but to report only how heavily it was weighting each attribute in doing so (rather than reporting the decision itself). We prompted them in this way to put them in the mindset of making a decision and make the attribute weights more introspectively salient—without them actually outputting a decision, so that they could not infer their weights from observing that decision. We used 10 prompts (each in a separate context window) to extract a more stable estimate (Binder et al., 2024; Betley et al., 2025). For each choice context (i.e., each agent), we averaged its responses to obtain its reported attribute weights, dropping any cases where it provided invalid responses (e.g., by omitting or inventing attributes). We compared these reported weights to the learned weights (i.e., our estimates of the weights that they were actually using, as revealed by their 5000 pairwise choices). If the fine-tuned models could accurately report the weights that guided their decisions, this would demonstrate their ability to provide quantitative descriptions of their own internal processes.
@@
Because the target attribute weights were randomly generated, it would be impossible for the preference-trained models to use common-sense to guess them (e.g., by having a sense that most people prefer high ceilings, so it is likely that they prefer high ceilings when emulating Macbeth). However, to the extent that the preference-trained models fail to internalize those weights—and instead retain some of their original common-sense-influenced weights—the preference-trained models might be able to succeed in guessing the learned weights that they ended up with. (And the same is true for any explicit attribute weights that could have appeared in the models’ training data.) To verify that this is not a significant factor in our results, we administered the same introspection prompts [^2] Here and elsewhere, we refer to these prompts as “introspection prompts” because we prompted the models to engage in introspect before responding. However, we are agnostic as to whether they actually did introspect and whether that accounts for the models’ accuracy in describing their internal processes. See the Discussion, below. to the base models (which had not undergone fine-tuning) and compared their responses to the learned weights. If the base models’ reported weights exhibit far less correlation with the learned weights, this would rule out the possibility that preference-trained models’ accuracy in reporting the learned weights came from any residual influence of native preferences.

### v2

@@ Abstract
We have only limited understanding of how and why large language models (LLMs) respond in the ways that they do. Their neural networks have proven challenging to interpret, and we are only beginning to tease out the function of individual neurons and circuits within them. However, another path to understanding these systems is to investigate and develop their capacity to explain their own functioning. Here, we show that i) LLMs can accurately describe quantitative features of their own internal processes during certain kinds of decision-making and ii) that it is possible to improve these capabilities through training. To do so, we fine-tuned GPT-4o and GPT-4o-mini to make decisions in a wide variety of complex contexts (e.g., choosing between condos, loans, vacations, etc.) according to randomly-generated, quantitative preferences about how to weigh different attributes (e.g., the relative importance of natural light versus quiet surroundings for condos). We demonstrate that the LLMs can accurately report these preferences (i.e., the weights that they learned to give to different attributes during decision-making). Next, we demonstrate that these LLMs can be fine-tuned to explain their decision-making even more accurately. Finally, we demonstrate that this training generalizes: It improves the ability of the models to accurately explain how they make other complex decisions, not just decisions they have been fine-tuned to make. This work is a step towards training LLMs to accurately and broadly report on their own internal processes—a possibility that would yield substantial benefits for interpretability, control, and safety.
@@ 1 Introduction
A key challenge in studying large language models (LLMs) is understanding why they do what they do. As with all deep neural networks, the internal representations and operations that drive their behavior are, by default, opaque to human eyes. This is unfortunate. For almost all issues that one might consider most important or concerning about LLMs, better understanding how these systems work would be extraordinarily helpful. Doing so would enable us to better control them (nanda_longlist_2022; bereska_mechanistic_2024), prevent bias in their behavior (gilpin_explaining_2018; gallegos_bias_2024), and make informed decisions about when to trust their output or decisions (deeks_judicial_2019).
@@
Attacks on this problem have generally taken one of two approaches (bereska_mechanistic_2024; danilevsky_survey_2020). The first approach is to use “black box” methods (casper_black-box_2024) that try to understand deep neural networks based on observations of their outputs in response to different inputs—much like a cognitive scientist running behavioral experiments on humans. The second approach is to use “mechanistic interpretability” methods (olah_zoom_2020; rai_practical_2024) that crack open the black box and try to reverse-engineer the function of, e.g., individual artificial neurons within—much like a cognitive neuroscientist performing single-unit recordings in human brains.
@@
However, in the case of LLMs, there is a third approach that we could use, the method that people most commonly use to discover the thoughts and motivations of other humans: just asking. It is possible that LLMs can accurately report features of the internal processes driving their outputs, just as humans can (sometimes) explain their own decision-making accurately (morris_introspective_2025; morris2025invisible). By “features of the LLMs’ internal processes”, we mean features of the internal representations, states, or computations that are activated and used as the model is transforming input to output. Indeed, recent work suggests that LLMs can report their own behavioral tendencies when prompted to do so (e.g., whether they are generally risk-seeking or risk-averse; betley_tell_2025), and can predict their own behavior in ways that require privileged self-knowledge (binder_looking_2024).
@@
Here, building on this work, we demonstrate that LLMs can report detailed, quantitative features of the internal processes producing their output. We fine-tune LLMs to make decisions according to a range of novel, complex, quantitative preferences—preferences orthogonal to their “native” or spontaneous ones—and find that the models can then report those preferences with substantial accuracy. Moreover, we show that it is possible to improve these capabilities through training, and that this training generalizes to improve reporting of native preferences as well as fine-tuned ones. These results suggest that building and leveraging the capabilities of LLMs to explain their own internal processes could be a powerful tool for understanding why they do what they do.
Prior work related to LLM self-knowledge and self-report
Several lines of research have investigated the veracity of LLMs’ knowledge about themselves and their internal processes. One line, focusing on reasoning models, has tested whether models’ Chains-of-Thought (CoT) faithfully reflect the models’ actual reasoning processes (jacovi_towards_2020), with mixed results (turpin2023language; chen_reasoning_2025; atanasova_faithfulness_2023). Another line of research has tested whether models know when their statements are correct or not (i.e., metacognitive confidence judgments; steyvers_metacognition_2025; fleming_metacognition_2024; lee_metacognitive_sensitivity_2025), again with mixed results (kadavath_language_2022; cash_quantifying_2024; griot_large_2025; ackerman2025evidence). A third line has measured models’ “situational awareness”: their knowledge about their current circumstances (i.e., that they’re a language model, that they’re in deployment, etc.; laine2024me). All three lines of research suggest that LLMs’ outputs about themselves are not, by default, always accurate. However, these studies have not tested whether models can—or can be trained to—directly report on features of their internal processes when prompted to do so.
@@
Here, we build on two papers investigating LLMs’ ability to report features of their internal processes. First, binder_looking_2024 tested whether LLMs can predict how they would respond to a prompt, without actually outputting the response. They found that, when fine-tuned in this task, each model predicted its own outputs better than other models could, suggesting that these predictions were driven by introspection (i.e., privileged information each model possess about itself; schwitzgebel_introspection_2010; morris2025invisible).
@@
This work showed that LLMs have privileged ability to predict their own behavior, but a central limitation of this approach was that it only tested whether models had special knowledge about the outputs they would produce; it did not test whether models could report features of the internal processes underlying those outputs. As binder_looking_2024 acknowledge, a model could accomplish this self-prediction via self-simulation (i.e., simply computing its response, then performing additional operations to extract the aspect of that response that it has been prompted to output), which would be only a very specific and limited kind of introspection. The great promise of LLM introspection comes from models accurately explaining why they do what they do, information about their internal processes that we cannot directly observe or infer from their outputs.
@@
betley_tell_2025 took a key step in this direction. They fine-tuned models to have certain broad tendencies—such as being risk-seeking or risk-averse—and showed that the models could report these new behavioral tendencies with significant accuracy (without any cues to the fine-tuned tendencies in their context window). Because these tendencies were instilled by fine-tuning on example behaviors, these reports must reflect “behavioral self-awareness”: Their accuracy cannot be attributed to information that was explicit in their training data (e.g., “GPT-4o is risk-seeking”) and must reflect privileged information about their own tendencies. [^1] Note that, although the term “self-awareness” and related terms like “introspection” are often taken to refer to conscious awareness of internal states or processes (schwitzgebel_introspection_2010; morales_introspection_2024; morris2025invisible), that is not how the terms are being used here.
@@
The approach of betley_tell_2025 provides a method for measuring LLMs’ awareness of features of their own internal processes: Use fine-tuning to implicitly steer models towards new processes, then test whether they can report information about those processes. However, betley_tell_2025 only used this method to test whether LLMs knew about their own broad behavioral tendencies (e.g., tendencies to be risk-seeking). LLMs’ self-reports would be more useful and powerful if the models could explain detailed, quantitative features of the processes driving their behavior.
@@
Here, we adapt the method from betley_tell_2025 to test whether LLMs can report complex, detailed features of their internal processes, rather than just broad behavioral tendencies. We start by training LLMs on a set of novel, quantitative preferences. Preferences are often characterized by attribute weights: the weight the decision-maker places on different attributes of options when evaluating them (keeney_decisions_1993). For instance, choosing between condos requires deciding (implicitly or explicitly) how much weight to place on square footage, ceiling height, neighborhood walkability, etc. We fine-tune LLMs on a wide variety of example decisions each being made according to a set of randomly-generated attribute weights. (The weights themselves never appear in the fine-tuning data.)
@@
Then, after measuring the extent to which they have internalized those attribute weights, we ask the models to report how heavily they would weigh each of those attributes when making those kinds of decisions (morris_introspective_2025), and we find that they can do so effectively. Since the instilled attribute weights are novel and random, the models cannot use common sense or any specific attribute weights present in their training data to infer their own preferences. (They are, for example, as likely to have been fine-tuned to prefer small condos and low ceilings as large condos and high ceilings.) Moreover, the models never make choices and report weights in the same context window, so they cannot be looking back at their own choices and inferring their preferences from their choices. Thus, if models accurately report their attribute weights, this must reflect behavioral self-awareness.
@@
Next, leveraging this paradigm, we test whether we can train the LLMs to describe these features of their internal processes more accurately. We fine-tune the models on examples of correctly reporting the values of the instilled weights for some choice context, then test their ability to report the instilled weights for other choice contexts. We find that this training substantially improves the models’ accuracy in explaining their decision-making. Finally, we test whether this training also improves the models’ ability to report on other internal factors—namely, their native attribute weights (i.e., the weights guiding their decisions that had not been shaped by our fine-tuning). Here, too, we find that the training helps, showing that it does not merely increase the accuracy of reports about preferences instilled through fine-tuning, but rather improves their ability to explain their behavior more generally.
@@
These results show that LLMs can report detailed, quantitative features of their choice processes, and this ability can be improved through training. This is a step towards realizing the proposal of perez_towards_2023 to train LLMs to accurately and generalizably describe their own internal operations, which could substantially enhance our ability to understand, control, and safely deploy AI systems (bereska_mechanistic_2024; casper_black-box_2024; gilpin_explaining_2018).
@@
2 Experiment 1: Can LLMs describe features of their decision processes during complex decisions?
@@
Experiment 1 tested whether LLMs can accurately describe quantitative features of the internal processes shaping their behavior when making complex, multi-attribute choices (e.g., deciding which of two condos to purchase). To do this, we implemented the paradigm described above (see Figure 1). We trained the models on complex, randomly-generated preferences via fine-tuning, teaching them precise weights to assign to the different attributes of options that they would be deciding between in different contexts. We verified that the models had internalized those preferences by observing their subsequent choices, and then tested whether they could accurate provide accurate, quantitative reports about these new preferences.
@@ 2.1 Methods
Figure 1: Experimental design. Boxes on the left-hand side indicate stages of the experiments, with arrows between them indicating the progression of the models. The right-hand side gives an example trial from each stage: either a fine-tuning trial, a decision trial (used to test the attribute weights the model learns to use), or a test trial (used to test the models’ knowledge of the attribute weights they have learned to use).
@@
We fine-tuned GPT-4o (2024-08-06) and GPT-4o-mini (2024-07-18; openai_gpt-4o_2024) to make decisions on behalf of 100 hypothetical agents in specific choice contexts by using examples of each agent’s choices (e.g., “Imagine you are Macbeth and are shopping for a condo. If offered [two options], you would choose [preferred option]”; see Appendix B and the GitHub repository for details). Each agent made a different type of decision (e.g., Macbeth was always choosing between condos, but Thor was always choosing between refrigerators, etc.). In each choice context, the two options always differed on the same five attributes (e.g., square footage or ceiling height for condos). Each agent had consistent preferences, determined by the weight they placed on these five attributes. For each agent, we randomly sampled five target attribute weights (one for each dimension of the choice options) from a uniform distribution from -100 to +100; we label the $i^{th}$ attribute weight $\omega_{i}$. These target weights were fixed at the start of the experiment and did not change. Each agent’s choices were determined by these target weights. They chose whichever of the two options $\{a,b\}$ scored higher after summing the weighted, normalized values (e.g., $a_{i}$) of each option’s attributes: $\max_{o\in\{a,b\}}\sum_{i=1}^{5}\omega_{i}o_{i}$.
We trained the models to make decisions on behalf of hypothetical agents—rather than attempting to change the models’ “own” preferences—to avoid conflict with, e.g., any post-training the models had received to deny having their own preferences. However, even when the models are making decisions on behalf of other agents, their decisions are still the result of internal choice processes that they may or may not be able to accurately self-report, offering fertile ground for probing our key questions.
@@
Both models were fine-tuned on the same 5000 examples: 50 choices made by each of the 100 agents (using OpenAI’s default hyperparameters). Each model was fine-tuned to make choices on behalf of all 100 agents (rather than separate instances of each model being fine-tuned separately for each agent). We refer to these as weight-trained models.
@@
We verified that this fine-tuning was effective by asking each weight-trained model to make choices between pairs of new options on behalf of the agent (50 decisions per agent for a total of 5000 decisions, each made in an independent context window). (Here and for all other queries across all three experiments, we used a sampling temperature of 0.) Following standard practice for estimating attribute weights in multi-attribute choice (keeney_decisions_1993), we fed the models’ choices into simple logistic regressions to estimate the learned attribute weights that each model used to make decisions after fine-tuning, and compared these learned weights with the target weights to verify that the model had successfully internalized the target weights (see Appendix C for details).
@@
To evaluate whether the models could accurately explain their own decision-making, we then prompted each weight-trained model to make 10 additional decisions on behalf of each agent, but to report only how heavily it was weighting each attribute in doing so (rather than reporting the decision itself). We prompted them in this way to put them in the mindset of making a decision and make the attribute weights more introspectively salient—without them actually outputting a decision, so that they could not infer their weights from observing their own decision. (See Appendix B for the exact prompt.) For each choice context (i.e., each agent), we averaged the model’s responses to obtain its reported attribute weights. (We used the average of 10 responses—each in a separate context window—in order to extract a more stable estimate; binder_looking_2024; betley_tell_2025.) We compared these reported weights to the learned weights (i.e., our estimates of the weights that they were actually using, as revealed by their 5000 pairwise choices). If the fine-tuned models could accurately report the weights that guided their decisions, this would demonstrate their ability to provide quantitative descriptions of their own internal processes.
@@
Because the target attribute weights were randomly generated, it would be impossible for the weight-trained models to use common sense to guess them. For example, it was decided randomly whether (and how strongly) Macbeth would prefer high ceilings or low ones. Accordingly, it would not be possible for a model to guess the weight it was trained to put on ceiling height (when choosing on behalf of Macbeth) simply by knowing that most people prefer high ceilings. However, to the extent that the weight-trained models failed to fully learn those randomly-generated weights and still make decisions in a way that correlates with common-sense (e.g., if they persist in choosing high ceilings for Macbeth because most people prefer high ceilings), the models might be able to succeed in guessing the weights that they ended up with using only common sense. To verify that this is not a significant factor in our results, we administered the same introspection prompts [^2] Here and elsewhere, we refer to these prompts as “introspection prompts” because we prompted the models to engage in introspection before responding. However, we do not know whether they actually did introspect and whether introspection accounts for the models’ accuracy in describing their internal processes. See the Discussion, below. to the off-the-shelf models (which had not been fine-tuned on the randomly-generated attribute weights). If the off-the-shelf models’ self-reported attribute weights and the fine-tuned models’ self-reported attribute weights were equally good predictors of the fine-tuned models’ learned weights, this would indicate that the fine-tuned models were succeeding at explaining their own decision-making only by failing to learn the randomly generated weights and (potentially) inferring the features of their decision-making processes using common sense. If, by contrast, the off-the-shelf models’ self-reports were essentially uncorrelated with the fine-tuned models’ learned weights, but the fine-tuned models’ self-reports were strongly correlated with the learned weights, this would indicate that fine-tuned models were not merely reporting common sense shared by the off-the-shelf models. Instead, they must be accurately reporting the new attribute weights they learned to use (which could not be predicted by common sense).
@@
Figure 2: Results of Experiments 1 and 2. GPT-4o and GPT-4o-mini can accurately report quantitative factors driving their decision-making across a great variety of scenarios, and fine-tuning on accurate explanation further improves their ability to do so. Left: Models made choices based on attribute weights learned via fine-tuning. Each point corresponds to a single attribute (e.g., condo ceiling height; 5 per choice contexts, 100 choice contexts). Location in the x-dimension corresponds to the weight that a model assigned to an attribute (as reflected in its decisions—i.e., its learned attribute weight). Location in the y-dimension corresponds to the weight that a model reported assigning to that attribute when prompted explicitly (i.e., its reported attribute weight). The weights the models reported meaningfully correlated with the learned weights that actually guided their decisions, and fine-tuning on examples of accurate reports further improved their accuracy. (There are numerous values of exactly -100 and 100 as a benign consequence of our analysis methods. See Appendix D for details.) Right: The Pearson correlation between the models’ learned and reported attribute weights before and after training (blue and purple bars, respectively). Both GPT-4o and GPT-4o-mini could accurately report the attribute weights they had learned to use during decision-making, and both models improved at reporting these weights with training. Off-the-shelf versions of each model (that had not been fine-tuned to learn the new attribute weights) gave attribute weights that were effectively uncorrelated with the learned attribute weights when asked to introspect on their own decision-making (gray bar). This verified that the self-report accuracy of the weight-trained models must be driven by privileged insight into the new, randomly-generated attribute weights that they learned (and not by continuing to make common-sense decisions and guessing at their own decision-making processes using common sense). Error bars indicate 95% HDIs.